OpenStudy (anonymous):

Gary has 2 apple pies and each is divided into 8 equal parts. He serves 4/8 of the first pie. How much pie does he have left?

OpenStudy (johnweldon1993):

okay so 2 pies....each in 8 parts...if he serves 4/8 of the first pie...he has..8/8-4/8 = 4/8 left of the first pie...but a whole second pie so 4/8 + 8/8 = 12/8 which equals 1 pie and 4/8 of another...of 1 and a half pies left

OpenStudy (anonymous):

thank you

OpenStudy (johnweldon1993):

no problem :)